OpenStudy (anonymous):

PLEASE HELP WITH RATIONAL EXPRESSION ??? CONSIDER THE FOLLOWING EQUATION -3/X-3=1 -2/X+7

OpenStudy (jdoe0001):

\(\bf \cfrac{-3}{x-3}=\cfrac{-2}{x+7}\implies \cfrac{?}{(x+7)(x-3)}\)

OpenStudy (jdoe0001):

hmmm.... wait... tis no addition.... so

OpenStudy (jdoe0001):

\(\bf \cfrac{-3}{x-3}=\cfrac{-2}{x+7}\quad \textit{let us cross-multiply}\\ \quad \\ \implies -3(x+7)=-2(x-3)\)

OpenStudy (jdoe0001):

solve for "x"

OpenStudy (anonymous):

\[\frac{-1}{x-3}=1-\frac{2}{x+7}\] like that?